# On finding optimal collective variables for complex systems by minimizing the deviation between effective and full dynamics

### v1

@@ 4.2 Transition rates
Next, we compare the transition rates of $Z_{n}$ to the corresponding transition rates of $X_{n}$. To begin with, let us recall that in TPT the transition rate $k_{AB}$ of $X_{n}$ given two disjoint closed subsets $A,B\subset\mathbb{R}^{d}$ is defined in (31) in Section 2.3. To make a connection to the effective dynamics, let us furthermore assume that these two subsets are defined using the CV map $\xi$, i.e. there are two disjoint subsets $\widetilde{A},\widetilde{B}\subset\mathbb{R}^{k}$, such that (see Figure 2 for illustration)
@@
$\widetilde{k}_{\widetilde{A}\widetilde{B}}=\widetilde{\mathcal{E}}(\widetilde{q})=\frac{1}{2}\int_{\mathbb{R}^{k}}\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})|\widetilde{q}(\hat{z})-\widetilde{q}(z)|^{2}Q(z)\mathrm{d}{z}\,\mathrm{d}{\hat{z}}\,,$ (76)
@@
We prove the claims by applying Proposition 4. Consider the function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ defined by $f=\widetilde{q}\circ\xi$. The conditions in (77b) and the relation (74) imply that (35) is satisfied by $f$. The equation (78) follows from the identity (36) in Proposition 4 and the fact that $\mathcal{E}(f)=\mathcal{E}(\widetilde{q}\circ\xi)=\widetilde{\mathcal{E}}(\widetilde{q})=\widetilde{k}_{\widetilde{A}\widetilde{B}}$, which in turn is implied by (65) and (76).
@@
We conclude this section with a discussion on learning the CV map $\xi$ by optimizing the transition rate estimation.
Remark 4.2 (Optimizing CV maps by transition rate estimation).
Similar to the discussions in Remark 4.1, we can also optimize the CV map $\xi$ (for the study of transition events) using the variational characterization in Proposition 4 established for the committor $q$. Specifically, we solve the problem (37) under the condition (35) among functions $f$ of the form $f=\widetilde{f}\circ\xi$, where $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ and $\widetilde{f}:\mathbb{R}^{k}\rightarrow\mathbb{R}$. Again, in this case Lemma 2 implies that the condition (35) on $f$ is equivalent to the condition (77b) on $\widetilde{f}$, and for the objective (37) we have
$\min_{f}\mathcal{E}(f)=\min_{\xi}\min_{\widetilde{f}}\mathcal{E}(\widetilde{f}\circ\xi)=\min_{\xi}\min_{\widetilde{f}}\widetilde{\mathcal{E}}(\widetilde{f})=\min_{\xi}\widetilde{k}_{\widetilde{A}\widetilde{B}}\,,$ (79)
where the last equality follows by applying Proposition 4 to the effective dynamics built with the map $\xi$. To summarize, this approach finds the optimal CV map $\xi$ such that the transition rate is best approximated by the corresponding effective dynamics.
@@ 4.3 Comparison under transformations
Finally, we compare timescales and transition rates associated to two effective dynamics. Besides the CV map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ and the corresponding effective dynamics $Z_{n}$, let us consider another CV map $\xi^{\prime}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k^{\prime}}$, which is related to $\xi$ by $\xi^{\prime}=f\circ\xi$, where $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{k^{\prime}}$ and $1\leq k^{\prime}\leq k<d$. This may correspond to the case discussed in Remark 3.3 where a CV map $\xi^{\prime}$ is built using system’s features (described by $\xi$).
@@
$\widetilde{p}\,^{\prime}(z^{\prime},\hat{z}^{\prime})=\frac{1}{Q^{\prime}(z^{\prime})}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi^{\prime}(x)-z^{\prime})\delta(\xi^{\prime}(y)-\hat{z}^{\prime})\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}\,,\quad z^{\prime},\hat{z}^{\prime}\in\mathbb{R}^{k^{\prime}}\,,$ (80)
where $Q^{\prime}(z^{\prime})=\int_{\mathbb{R}^{d}}\delta(\xi^{\prime}(x)-z^{\prime})\pi(x)\,\mathrm{d}{x}$ is a normalizing factor. Since $X_{n}$ is assumed to be reversible, Corollary 1 implies that $Z_{n}^{\prime}$ is reversible with respect to its invariant probability measure $\widetilde{\mu}^{\prime}$, where $\mathrm{d}{\widetilde{\mu}^{\prime}}=Q^{\prime}(z^{\prime})\mathrm{d}{z^{\prime}}$. Let $\widetilde{\mathcal{T}}^{\prime}$ be the corresponding transfer operator of $Z_{n}^{\prime}$, with (nonnegative) eigenvalues $1=\widetilde{\lambda}^{\prime}_{0}>\widetilde{\lambda}^{\prime}_{1}\geq\widetilde{\lambda}^{\prime}_{2}\geq\dots$, and eigenfunctions $(\widetilde{\varphi}^{\prime}_{i})_{i\geq 0}$, which are normalized with respect to the norm $\|\cdot\|_{\widetilde{\mu}^{\prime}}$. Given the subsets $\widetilde{A},\widetilde{B}\subset\mathbb{R}^{k}$ that were considered in the previous section (see (74)), we assume that there are two subsets $\widetilde{A}^{\prime},\widetilde{B}^{\prime}\subset\mathbb{R}^{k^{\prime}}$, such that $\widetilde{A}=f^{-1}(\widetilde{A}^{\prime})$, $\widetilde{B}=f^{-1}(\widetilde{B}^{\prime})$ and $(\widetilde{A}^{\prime}\cup\widetilde{B}^{\prime})^{c}\neq\emptyset$. Let $\widetilde{k}^{\prime}_{\widetilde{A}^{\prime}\widetilde{B}^{\prime}}$ denote the transition rate of $Z_{n}^{\prime}$ from the set $\widetilde{A}^{\prime}$ to the set $\widetilde{B}^{\prime}$.
We have the following result, which elucidates the relation between the effective dynamics $Z_{n}$ and $Z_{n}^{\prime}$ (see Figure 3), and also the relations between their timescales and transition rates.
@@
Concerning the first claim, for the integral on the right hand side of (80) we can derive
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi^{\prime}(x)-z^{\prime})\delta(\xi^{\prime}(y)-\hat{z}^{\prime})\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$
$=$ $\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta\big{(}f(\xi(x))-z^{\prime}\big{)}\delta\big{(}f(\xi(y))-\hat{z}^{\prime}\big{)}\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$
$=$ $\int_{\mathbb{R}^{k}}\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi(x)-z)\delta(\xi(y)-\hat{z})\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}\bigg{]}$
$\hskip 42.67912pt\delta(f(z)-z^{\prime})\delta(f(\hat{z})-\hat{z}^{\prime})\,\mathrm{d}{z}\,\mathrm{d}{\hat{z}}$
$=$ $\int_{\mathbb{R}^{k}}\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\delta(f(z)-z^{\prime})\delta(f(\hat{z})-\hat{z}^{\prime})Q(z)\,\mathrm{d}{z}\,\mathrm{d}{\hat{z}}\,,$
where the first equality follows from the relation $\xi^{\prime}=f\circ\xi$, the second equality follows by applying the identity (42), and the last equality follows from the definition of $\widetilde{p}$ in (44). Similarly, for the normalizing factor on the right hand side of (80) we have
$Q^{\prime}(z^{\prime})=$ $\int_{\mathbb{R}^{d}}\delta\big{(}f(\xi(x))-z^{\prime}\big{)}\pi(x)\,\mathrm{d}{x}$
$=$ $\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{d}}\delta(\xi(x)-z)\pi(x)\,\mathrm{d}{x}\bigg{]}\delta(f(z)-z^{\prime})\,\mathrm{d}{z}$
$=$ $\int_{\mathbb{R}^{k}}\delta(f(z)-z^{\prime})Q(z)\,\mathrm{d}{z}\,,$
where the last equality follows from (40). Substituting the above expressions into (80), noticing the fact that $\widetilde{p}$ and $Q(z)$ are the transition density and the invariant density of $Z_{n}$ respectively, and using the definition of effective dynamics in Section 3 (see (40) and Definition 44), we conclude that $Z_{n}^{\prime}$ is the effective dynamics of $Z_{n}$ associated to the CV map $f$.
@@
As an application of Proposition 8, we obtain the following result when $k^{\prime}=k$ and $f$ is an isomorphism (bijection).
Corollary 2.
Assume that $k^{\prime}=k$ and $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{k}$ is an isomorphism. Then the effective dynamics $Z_{n}^{\prime}$ is identical to $Z_{n}$ up to the transformation $f$. In particular, both processes $Z_{n}$ and $Z_{n}^{\prime}$ have the same eigenvalues and transition rates.
Proof.
Let us consider two test functions $\widetilde{g}^{\prime}_{1},\widetilde{g}^{\prime}_{2}:\mathbb{R}^{k^{\prime}}\rightarrow\mathbb{R}$. Using the fact that $f$ is an isomorphism, the identity (42), and the expressions (44),(80), we can derive
$\int_{\mathbb{R}^{k^{\prime}}}\int_{\mathbb{R}^{k^{\prime}}}\widetilde{p}^{\prime}(z^{\prime},\hat{z}^{\prime})\widetilde{g}^{\prime}_{1}(z^{\prime})\widetilde{g}^{\prime}_{2}(\hat{z}^{\prime})Q^{\prime}(z^{\prime})\,\mathrm{d}{z}^{\prime}\mathrm{d}{\hat{z}}^{\prime}$
$=$ $\int_{\mathbb{R}^{k^{\prime}}}\int_{\mathbb{R}^{k^{\prime}}}\bigg{[}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi^{\prime}(x)-z^{\prime})\delta(\xi^{\prime}(y)-\hat{z}^{\prime})\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}\bigg{]}\widetilde{g}^{\prime}_{1}(z^{\prime})\widetilde{g}^{\prime}_{2}(\hat{z}^{\prime})\,\mathrm{d}{z}^{\prime}\mathrm{d}{\hat{z}}^{\prime}$
$=$ $\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\widetilde{g}^{\prime}_{1}(f(\xi(x)))\widetilde{g}^{\prime}_{2}(f(\xi(y)))\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$
$=$ $\int_{\mathbb{R}^{k}}\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\widetilde{g}^{\prime}_{1}(f(z))\widetilde{g}^{\prime}_{2}(f(\hat{z}))Q(z)\,\mathrm{d}{z}\mathrm{d}{\hat{z}}\,.$
Since the test functions $\widetilde{g}^{\prime}_{1},\widetilde{g}^{\prime}_{2}$ are arbitrary, we can obtain the relations (choosing $\widetilde{g}^{\prime}_{2}\equiv 1$ first)
$Q^{\prime}(z^{\prime})\mathrm{d}{z}^{\prime}=Q(z)\mathrm{d}{z},\quad\widetilde{p}\,^{\prime}(z^{\prime},\hat{z}^{\prime})\,\mathrm{d}{\hat{z}}^{\prime}=\widetilde{p}(z,\hat{z})\,\mathrm{d}{\hat{z}}\,,$
for all $z^{\prime},\hat{z}^{\prime}\in\mathbb{R}^{k^{\prime}}$, where $z=f^{-1}(z^{\prime}),\hat{z}=f^{-1}(\hat{z}^{\prime})$. This shows that the processes $Z_{n}$ and $Z_{n}^{\prime}$ are identical up to the transformation $f$. ∎
We conclude this section with a discussion on the practical implication of Proposition 8.
Remark 4.3.
Proposition 8 relates different steps of a multi-step model reduction procedure in terms of effective dynamics and it shows that the approximation errors aggregate in each step. Together with Theorem 4.1 and Theorem 4.3, it also allows to inspect approximation errors (layer-by-layer) in deep learning approaches where multi-layer artificial neural networks are employed to represent CV maps.
@@ 5 Conclusion and Discussions
There are manifold connections between the work presented herein and similar approaches proposed in the literature. For example, the effective dynamics of diffusion processes described by SDEs have been studied in several works and different types of error estimates were obtained, e.g., in [25, 54, 26, 8, 30], where the error estimates obtained in this work are deeply related to the estimates in [54]. Beyond this, the quality of effective dynamics in approximating the original process can also be studied by considering the relative entropy or pathwise error between the effective dynamics and the original process, as is done, e.g., in [25, 26, 8, 30] for diffusion processes.
While the main results of this work are theoretical, we discussed the relevance of effective dynamics in data-driven numerical approaches, such as in VAMPnets [31, 6], deep learning of transition manifolds [3] and existing algorithms for learning CV maps [51, 27]. The approach taken herein, however, may open the horizon for algorithmic approaches for learning ”good” CVs and the associated effective dynamics simultaneously where the already learned effective dynamics might also be utilized for exploring state space, a key step that is still the main bottleneck of available algorithmic approaches. In future work, this approach will be further studied and applied to concrete systems in MD applications.
@@ Acknowledgments
CS is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy MATH+: The Berlin Mathematics Research Centre (EXC-2046/1, project No. 390685689) and CRC 1114 “Scaling Cascades in Complex Systems” (project No. 235221301). WZ is funded by DFG’s Eigene Stelle (project No. 524086759). WZ thanks Tony Lelièvre and Gabriel Stolz for inspiring discussions on the topic regarding collective variables.
@@ Appendix A Transfer operator for Langevin dynamics
Suppose that we want to construct a Markov chain for the data $(X_{n})_{0\leq n\leq N}$, where $X_{n}=x(n\tau)$ is the position of (17) at time $t=n\tau$ and $\tau>0$ is the lag-time. Let $p(x,y)$ be the transition density of the Markov chain estimated from the data when the length $N\rightarrow+\infty$. Given $\bar{x},\bar{y}\in\mathbb{R}^{d}$, let us denote by $D_{1},D_{2}$ small neighborhoods of $\bar{x}$ and $\bar{y}$, respectively. We can derive
$\int_{D_{2}}p(\bar{x},y)\,\mathrm{d}{y}$
$=$ $\lim_{|D_{1}|\rightarrow 0}\frac{1}{|D_{1}|}\int_{D_{2}}\int_{D_{1}}p(x,y)\,\mathrm{d}{x}\,\mathrm{d}{y}$
@@
$=$ $\big{(}\frac{\beta}{2\pi}\big{)}^{\frac{d}{2}}\lim_{|D_{1}|\rightarrow 0}\frac{\int_{D_{2}}\int_{D_{1}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{\mathrm{Lan}}(\tau,y,v^{\prime}|x,v)\,\mathrm{e}^{-\beta(V(x)+\frac{|v|^{2}}{2})}\,\mathrm{d}{v}\,\mathrm{d}{v^{\prime}}\,\mathrm{d}{x}\,\mathrm{d}{y}}{\int_{D_{1}}\mathrm{e}^{-\beta V(x)}\,\mathrm{d}{x}}$
$=$ $\big{(}\frac{\beta}{2\pi}\big{)}^{\frac{d}{2}}\int_{D_{2}}\Big{[}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{\mathrm{Lan}}(\tau,y,v^{\prime}|\bar{x},v)\,\mathrm{e}^{-\frac{\beta|v|^{2}}{2}}\,\mathrm{d}{v}\,\mathrm{d}{v^{\prime}}\Big{]}\,\mathrm{d}{y}\,,$
where the first and the last equalities follow from the continuity of densities $p$ and $p^{\mathrm{Lan}}$, respectively, the second equality follows from the definition of $p$ (i.e. the way in which $p$ is estimated from data), the third equality follows from the ergodicity of the process. Since the above derivation is true for any neighborhood $D_{2}$ of $\bar{y}$, we conclude that
$p(\bar{x},\bar{y})=\big{(}\frac{\beta}{2\pi}\big{)}^{\frac{d}{2}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{\mathrm{Lan}}(\tau,\bar{y},v^{\prime}|\bar{x},v)\,\mathrm{e}^{-\frac{\beta|v|^{2}}{2}}\,\mathrm{d}{v}\,\mathrm{d}{v^{\prime}}\,.$ (82)

### v2

@@ 4.2 Transition rates
Next, we compare the transition rates of $Z_{n}$ to the corresponding transition rates of $X_{n}$. To begin with, recall that in TPT the transition rate $k_{AB}$ of $X_{n}$ given two disjoint closed subsets $A,B\subset\mathbb{R}^{d}$ is defined in (30) in Section 2.3. To make a connection to the effective dynamics, let us furthermore assume that these two subsets are defined using the CV map $\xi$, i.e. there are two disjoint subsets $\widetilde{A},\widetilde{B}\subset\mathbb{R}^{k}$, such that (see Figure 2 for illustration)
@@
$\widetilde{k}_{\widetilde{A}\widetilde{B}}=\widetilde{\mathcal{E}}(\widetilde{q})=\frac{1}{2}\int_{\mathbb{R}^{k}}\left(\int_{\mathbb{R}^{k}}\widetilde{p}(z,w)(\widetilde{q}(w)-\widetilde{q}(z))^{2}\mathrm{d}{w}\right)\widetilde{\mu}(\mathrm{d}{z})\,,$ (79)
@@
We prove the claims by applying Proposition 4. Consider the function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ defined by $f=\widetilde{q}\circ\xi$. The conditions in (80b) and the relation (77) imply that $f\in\mathcal{F}_{AB}$ (see (33)). The equation (81) follows from the identity (34) in Proposition 4 and the fact that $\mathcal{E}(f)=\mathcal{E}(\widetilde{q}\circ\xi)=\widetilde{\mathcal{E}}(\widetilde{q})=\widetilde{k}_{\widetilde{A}\widetilde{B}}$, which in turn is implied by (68) and (79).
@@ 4.3 Comparison under transformations
Finally, we compare timescales and transition rates for effective dynamics associated to two CV maps. Besides the CV map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ and the corresponding effective dynamics $Z_{n}$, let us consider another CV map $\xi^{\prime}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k^{\prime}}$, which is related to $\xi$ by $\xi^{\prime}=f\circ\xi$, where $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{k^{\prime}}$ and $1\leq k^{\prime}\leq k<d$. This may correspond to the case discussed in Remark 3.2 where a CV map $\xi^{\prime}$ is built using system’s features (described by $\xi$).
@@
$\widetilde{p}\,^{\prime}(z^{\prime},w^{\prime})=\int_{\Sigma_{z^{\prime}}}\Big{[}\int_{\Sigma_{w^{\prime}}}p(x,y)[\mathrm{det}(\nabla\xi^{\prime}(y)^{\top}\nabla\xi^{\prime}(y))]^{-\frac{1}{2}}\sigma_{\Sigma_{w^{\prime}}}(\mathrm{d}{y})\Big{]}\,\mu_{z^{\prime}}(\mathrm{d}{x})\,,$ (82)
where $z^{\prime},w^{\prime}\in\mathbb{R}^{k^{\prime}}$. Since we assume $X_{n}$ to be reversible, Corollary 1 implies that $Z_{n}^{\prime}$ is reversible with respect to its invariant probability measure $\widetilde{\mu}^{\prime}(\mathrm{d}{z^{\prime}})=\widetilde{\pi}^{\prime}(z^{\prime})\mathrm{d}{z^{\prime}}$, where $\widetilde{\pi}^{\prime}(z^{\prime})=\int_{\Sigma_{z^{\prime}}}\pi(x)[\mathrm{det}(\nabla\xi^{\prime}(x)^{\top}\nabla\xi^{\prime}(x))]^{-\frac{1}{2}}\,\sigma_{\Sigma_{z^{\prime}}}(\mathrm{d}{x})$ (see (39)). Let $\widetilde{\mathcal{T}}^{\prime}$ be the corresponding transfer operator of $Z_{n}^{\prime}$, with (nonnegative) eigenvalues $1=\widetilde{\lambda}^{\prime}_{0}>\widetilde{\lambda}^{\prime}_{1}\geq\widetilde{\lambda}^{\prime}_{2}\geq\dots$, and eigenfunctions $(\widetilde{\varphi}^{\prime}_{i})_{i\geq 0}$, which are normalized with respect to the norm $\|\cdot\|_{\widetilde{\mu}^{\prime}}$. Given the subsets $\widetilde{A},\widetilde{B}\subset\mathbb{R}^{k}$ that were considered in the previous section (see (77)), we assume that there are two subsets $\widetilde{A}^{\prime},\widetilde{B}^{\prime}\subset\mathbb{R}^{k^{\prime}}$, such that $\widetilde{A}=f^{-1}(\widetilde{A}^{\prime})$, $\widetilde{B}=f^{-1}(\widetilde{B}^{\prime})$ and $(\widetilde{A}^{\prime}\cup\widetilde{B}^{\prime})^{c}\neq\emptyset$. Let $\widetilde{k}^{\prime}_{\widetilde{A}^{\prime}\widetilde{B}^{\prime}}$ denote the transition rate of $Z_{n}^{\prime}$ from the set $\widetilde{A}^{\prime}$ to the set $\widetilde{B}^{\prime}$.
The following result elucidates the relation between the effective dynamics $Z_{n}$ and $Z_{n}^{\prime}$ (see Figure 3), and also the relations between their timescales and transition rates.
@@
Concerning the first claim, for two test functions $\widetilde{h}^{\prime}_{1},\widetilde{h}^{\prime}_{2}:\mathbb{R}^{k^{\prime}}\rightarrow\mathbb{R}$, using (82) and the formula (40) we can derive
$\int_{\mathbb{R}^{k^{\prime}}}\left(\int_{\mathbb{R}^{k^{\prime}}}\widetilde{p}\,^{\prime}(z^{\prime},w^{\prime})\widetilde{h}^{\prime}_{1}(w^{\prime})\mathrm{d}{w^{\prime}}\right)\widetilde{h}^{\prime}_{2}(z^{\prime})\widetilde{\mu}^{\prime}(\mathrm{d}{z^{\prime}})$ (83)
$=$ $\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}p(x,y)\widetilde{h}_{1}^{\prime}(\xi^{\prime}(y))\mathrm{d}{y}\right)\widetilde{h}_{2}^{\prime}(\xi^{\prime}(x))\mu(x)$
$=$ $\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}p(x,y)\widetilde{h}_{1}^{\prime}(f(\xi(y)))\mathrm{d}{y}\right)\widetilde{h}_{2}^{\prime}(f(\xi(x)))\mu(x)$
$=$ $\int_{\mathbb{R}^{k}}\left(\int_{\mathbb{R}^{k}}\widetilde{p}(z,w)\widetilde{h}^{\prime}_{1}(f(w))\mathrm{d}{w}\right)\widetilde{h}^{\prime}_{2}(f(z))\widetilde{\mu}(\mathrm{d}{z})\,.$
In particular, taking $\widetilde{h}^{\prime}_{1}\equiv 1$, the above identities simplify to
$\int_{\mathbb{R}^{k^{\prime}}}\widetilde{h}^{\prime}_{2}(z^{\prime})\widetilde{\mu}^{\prime}(\mathrm{d}{z^{\prime}})=\int_{\mathbb{R}^{k}}\widetilde{h}^{\prime}_{2}(f(z))\widetilde{\mu}(\mathrm{d}{z})\,,$ (84)
which implies that $\widetilde{\mu}^{\prime}$ is the pushforward measure of $\widetilde{\mu}$ under $f$. Notice that $\widetilde{p}$ and $\widetilde{\mu}$ are the transition density and the invariant measure of $Z_{n}$ respectively. Using the definition of effective dynamics (see (39) and Definition 43) and rewriting the last expression in (83) using co-area formula [37, Lemma 3.2], we conclude that $Z_{n}^{\prime}$ is the effective dynamics of $Z_{n}$ associated to the CV map $f$.
@@ 5 Conclusion and Discussions
There are manifold connections between the work presented herein and similar approaches proposed in the literature. For example, the effective dynamics of diffusion processes described by SDEs have been studied in several works and different types of error estimates were obtained, e.g., in [34, 66, 35, 14, 39]. The error estimates in Section 4 of this work are closely related to the estimates in [66]. Beyond this, the quality of effective dynamics in approximating the original process can also be studied by considering the relative entropy or pathwise error between the effective dynamics and the original process, as is done, e.g., in [34, 35, 14, 39] for diffusion processes. However, to our best knowledge, there are no similar results on the optimal characterization for effective dynamics of diffusion processes in terms of relative entropy.
While the main results of this work are theoretical, we discussed the relevance of effective dynamics in data-driven numerical approaches, such as in VAMPnets [40, 12], deep learning of transition manifolds [7] and existing algorithms for learning CV maps [62, 36]. The approach taken herein, however, may open the horizon for algorithmic approaches for learning “good” CVs and the associated effective dynamics simultaneously where the already learned effective dynamics might also be utilized for exploring state space, a key step that is still the main bottleneck of available algorithmic approaches. In future work, this approach will be further studied and applied to concrete systems in MD applications.
@@ Acknowledgments
CS is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy MATH+: The Berlin Mathematics Research Centre (EXC-2046/1, project No. 390685689) and CRC 1114 “Scaling Cascades in Complex Systems” (project No. 235221301). WZ is funded by DFG’s Eigene Stelle (project No. 524086759). WZ thanks Tony Lelièvre and Gabriel Stolz for inspiring discussions on topics related to collective variables.
@@ Appendix A Transfer operator for Langevin dynamics
Suppose that we want to construct a Markov chain for the data $(X_{n})_{0\leq n\leq N}$, where $X_{n}=x(n\tau)$ is the position of (17) at time $t=n\tau$ and $\tau>0$ is the lag-time. Let $p(x,y)$ be the transition density of the Markov chain estimated from the data when the length $N\rightarrow+\infty$. Given $x,y\in\mathbb{R}^{d}$, let us denote by $D_{1},D_{2}$ small neighborhoods of $x$ and $y$, respectively. We can derive
$\int_{D_{2}}p(x,\bar{y})\,\mathrm{d}{\bar{y}}$
$=$ $\lim_{|D_{1}|\rightarrow 0}\frac{1}{|D_{1}|}\int_{D_{2}}\int_{D_{1}}p(\bar{x},\bar{y})\,\mathrm{d}{\bar{x}}\,\mathrm{d}{\bar{y}}$
@@
$=$ $\big{(}\frac{\beta}{2\pi}\big{)}^{\frac{d}{2}}\lim_{|D_{1}|\rightarrow 0}\frac{\int_{D_{2}}\int_{D_{1}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{\mathrm{Lan}}(\tau,\bar{y},v^{\prime}|\bar{x},v)\,\mathrm{e}^{-\beta(V(\bar{x})+\frac{|v|^{2}}{2})}\,\mathrm{d}{v}\,\mathrm{d}{v^{\prime}}\,\mathrm{d}{\bar{x}}\,\mathrm{d}{\bar{y}}}{\int_{D_{1}}\mathrm{e}^{-\beta V(\bar{x})}\,\mathrm{d}{\bar{x}}}$
$=$ $\big{(}\frac{\beta}{2\pi}\big{)}^{\frac{d}{2}}\int_{D_{2}}\Big{[}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{\mathrm{Lan}}(\tau,\bar{y},v^{\prime}|x,v)\,\mathrm{e}^{-\frac{\beta|v|^{2}}{2}}\,\mathrm{d}{v}\,\mathrm{d}{v^{\prime}}\Big{]}\,\mathrm{d}{\bar{y}}\,,$
where the first and the last equalities follow from the continuity of densities $p$ and $p^{\mathrm{Lan}}$, respectively, the second equality follows from the definition of $p$ (i.e. the way in which $p$ is estimated from data), the third equality follows from Birkhoff’s ergodic theorem (see [31, Theorem 4.4 in Chapter 1]) and the ergodicity of the process. Since the above derivation is true for any neighborhood $D_{2}$ of $y$, we conclude that
$p(x,y)=\big{(}\frac{\beta}{2\pi}\big{)}^{\frac{d}{2}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{\mathrm{Lan}}(\tau,y,v^{\prime}|x,v)\,\mathrm{e}^{-\frac{\beta|v|^{2}}{2}}\,\mathrm{d}{v}\,\mathrm{d}{v^{\prime}}\,.$ (86)